# HAMLOCK: HArdware-Model LOgically Combined attacK

### v1

@@ 4 Experiment Setup
We describe the experimental setup in Section˜4.1 and then introduce model-level defenses for evaluation in Section˜4.2.
@@ 4.1 Experimental Setup
Backdoor related settings. The backdoor triggers on these images are $3\times 3$ squares whose values are either fixed (single neuron weight optimization and multi-neuron attacks) or optimized based on Eq. (2) (single neuron trigger optimization attack). These patch-based triggers are the most primitive and easy to detect triggers for model-level attacks, as demonstrated in prior works [61, 16, 25], and we will then show that hardware model co-attack can make these highly detectable triggers highly evasive. We expect more sophisticated triggers will further improve the stealthiness of our attack. Some illustrative examples of trigger patterns for different datasets are shown in Figure˜5 in the Appendix. For our multi-neuron attack, we randomly select 3 layers, each with 1 trigger neuron.
@@
For model-level detections, we report the number of times our backdoored model is flagged as malicious, consistent with prior work [4]. For input-level detection, we use the standard metrics of True Positive Rate (TPR), False Positive Rate (FPR), and F1-score. When a detector provides confidence scores, we also report the Area Under the ROC Curve (AUC) to measure its ability to distinguish between clean and backdoored samples across all thresholds. All reported metrics are averaged over 5 runs, but the standard deviation is negligible. Our source code is publically available at https://github.com/Imsanskar/HAMLOCK.
@@ 4.2 Model-level Defenses
We select representative state-of-the-art model-level defenses to assess the stealth of HAMLOCK. These defenses are grouped into four categories: (1) Backdoor model detection: Neural Cleanse [61], MNTD [65]; (2) White-box backdoor sample detection: IBD-PSC [25] and TED [41]; (3) Black-box backdoor sample detection: STRIP [16] and BBCAL [26]; and (4) Backdoor mitigation: Finetuning [49], FinePruning [36] and BEAGLE [9]. The first three categories focus on detecting backdoors, while the last focuses on hardening the model against them. For more information about the defenses, we refer the reader to Appendix˜B. Our attack is a post-training supply-chain attack [4] and hence, is not compatible with during-training defenses [56, 21].
@@ 5.1 Effectiveness without Defense
First, it maintains a high clean accuracy (CA). Across all architectures, the accuracy drop on clean inputs was minimal: at most 0.3% for MNIST, 1.7% for CIFAR-10, 1.0% for GTSRB, and 2.6% for ImageNet. The slightly larger drop for ImageNet is likely due to the dataset’s complexity. Second, the model’s attack success rate (ASR) is effectively zero. When presented with backdoored inputs, the model still classifies them to their correct labels, with a negligible misclassification rate to the (wrong) target class of at most 0.6%. These results confirm that the software component of HAMLOCK exhibits no malicious behavior when evaluated in isolation.
However, once the model is deployed on the trojaned hardware, the attack becomes fully active. The hardware Trojan detects the trigger and activates its payload, causing the ASR to jump to 100% across all datasets and architectures. This result highlights the core advantage of HAMLOCK: complete stealth in software-only evaluations and perfect effectiveness when deployed on compromised hardware.
@@ 5.2 Evading White-box Defenses
This failure stems from a violation of each defense’s core assumptions. NC is ineffective because it searches for a small trigger pattern that causes misclassification. Our backdoored model, despite using a trigger that is very easy to reverse engineering with NC, it never misclassifies triggered inputs on its own, leaving NC with no trigger-label connection to find. Similarly, MNTD fails because it looks for statistical anomalies in neuron weights, but our attack’s minimal modification to few neurons does not create a detectable anomaly. This demonstrates that defenses examining a model’s behavior or structure are ill-equipped for a co-design attack where the malicious logic remains dormant in the software.
@@
Backdoor sample detection. We next evaluate our models against defences that detect backdoor samples, and results are in Table˜2. These defenses are also concurrent or prior to the release of DFBA but are not included in the original paper. Across all experimental settings, our three attack variants consistently evade detection. This is evidenced by uniformly low TPR (at most 12.4%), FPR (at most 11.5%), and F1 scores (at most 0.2). The proximity between TPR and FPR, along with AUC scores near 0.5, confirms that these defenses perform no better than random guessing at distinguishing our backdoored samples from clean ones.
@@
Effectiveness under Lightweight Retraining. Finally, we evaluate HAMLOCK against retraining defenses like fine-tuning and fine-pruning, with results presented in Table˜5. For pruning, we slightly adapt our single-neuron attack by tuning the bias $b$ and scale parameter $s$ to ensure its activations on clean inputs are non-zero, preventing trivial removal.
Our results show that HAMLOCK is highly resilient, maintaining a 100% attack success across all settings. Notably, this includes resilience against the advanced Beagle fine-tuning strategy, under a worst-case assumption where the defender has access to backdoor samples generated exactly from our attack. The original Beagle defense often analyzes backdoor samples from other proxy attacks. This resilience is twofold. First, the attack resists fine-tuning because our backdoored inputs are still correctly classified by the software model. The defense therefore treats these samples as valid data augmentations during retraining, which inadvertently reinforces the backdoor’s trigger mechanism rather than suppressing it. Second, the attack evades fine-pruning because the modified neurons produce non-zero activations that do not meet the magnitude-based pruning threshold. This demonstrates a strong resistance of our attack to common model-hardening techniques.
@@ 5.3 Evading Black-box Defenses
For completeness, we also evaluate black-box defenses that only require input-output access to the model. We test the robustness of both HAMLOCK and DFBA attacks against two defenses: the classic STRIP [16] and the more recent BBCAL [26]. Crucially, for HAMLOCK, we evaluate two distinct and realistic scenarios: 1) the dormant software model, as would be inspected by a Model Zoo maintainer before deployment, and 2) the active, hardware-hosted model, as would be tested by a cautious end-user. The latter case is unique because the hardware enables misclassifications on backdoored inputs while the software-only model does not.
The results of MNIST, GTSRB, CIFAR10, and Imagenet are presented in Table˜3 and Table˜4. Our HAMLOCK attack successfully bypasses both defenses in both scenarios. While BBCAL sometimes shows a high TPR, this comes at the cost of an equally high FPR, resulting in AUC scores near 0.5—no better than random guessing. The baseline DFBA attack is also largely evasive, with one minor exception: on ImageNet, STRIP achieves a 0.63 AUC score, which, however, is still considered ineffective for reliable detection.
@@ 5.4.1 Negligible Hardware Overhead
The hardware footprint is usually measured by additional area and power added to the overall circuit. A significant area and/or power overhead can be detected using side-channel analysis [54, 30, 5], while small overheads will simply be masked by the variation of the hardware costs during the fabrication process, making the detection of these footprints almost impossible. Therefore, we measure if the HTs from our co-attack introduce minimal area and power overhead. Table˜6 summarizes the results for the three model architectures considered in this paper: LeNet, VGG-16 and ResNet-18. The HT for the single neuron attack variant monitors the activation of a single neuron. The area overhead is capped at 0.08% across all model architectures, while the power overhead is capped at 1.14% for the VGG-16 model, and the rest simply drops to 0.02%. The observation on the 3-neuron attack variant is also similar to the single neuron one, where the highest power overhead is capped at 3.4% for VGG16, while the rest are similarly below 0.05%. The area overhead is negligible with the maximum overhead being 0.1%.
@@ 5.4.2 Diverse Trigger Conditions
Our hardware-based backdoor approach allows us to compose multiple, simple triggers into diverse and sophisticated activation conditions, a capability exceeding that of software-only attacks. We demonstrate three such compositions: combinational, sequential, and temporal triggers, all can be implemented with negligible hardware overhead.
A combinational trigger uses simple hardware logic (e.g., AND/OR gates) to combine the outputs of multiple, independent trigger detectors. For example, an AND gate requires multiple conditions to be met simultaneously (e.g., a specific road sign and foggy weather in an autonomous driving), while an OR gate allows any one of several conditions to activate the payload. Such a strategy allows attacker to split a complex trigger into smaller, stealthier pieces that are only malicious when they co-occur. We perform a preliminary experiment using two individual triggers, $3\times 3$ triggers at the bottom right and bottom left corners and each of them individually trigger a unique neuron, using the weight optimization technique in Section˜3.2.1. The results are shown in Table˜7, where the AND and OR logic are implemented with negligible overhead.
Beyond simple combinations, the hardware enables finite state machine (FSM) based complex sequential triggers. A sequential trigger activates the payload only after multiple, distinct trigger patterns have been observed in a specific order over time. A temporal trigger, implemented with a simple counter circuit, activates only after a set duration has passed or a certain number of inferences have been made. This allows for long-term dormant attacks, such as a backdoor in an autonomous vehicle that only manifests after a specific mileage, making the resulting failure indistinguishable from natural system degradation.
@@ 6 Related Work
Model-level Backdoor Attacks. Existing backdoor attacks include data poisoning [18, 57, 47, 7], model parameter modification [4, 23], and model architecture modification [3, 52] approaches. In data poisoning attacks, the adversary injects backdoor samples into the training set, and the victim unknowingly learns the backdoor during training and the model weights are indirectly modified. Other backdoors attacks alter the model weights directly [23, 4] or make architectural changes [3, 45]. Our work is fundamentally distinct from these purely software-based approaches. The HAMLOCK software model contains no functional backdoor path and does not cause misclassifications on its own, making it significantly stealthier, as demonstrated in Section˜5.2.
@@
Model-level Backdoor Defenses. Existing model-only defenses include during-training and post-training defenses. During-training defenses require filtering our bad training data [56, 29, 35] or suppress negative impact from the bad data [63, 53, 27]. Since our attack is a supply-chain attack [4, 18] introduced post training, these defenses do not apply. Post-training defenses, shown in Section˜5.2 to be ineffective against our attack, can be grouped into three categories. First are backdoored model detection methods, which aim to identify if a model has been compromised [61, 62, 64]. Second, model hardening defenses attempt to remove the backdoor’s effect through techniques like lightweight retraining [49, 36], adversarial unlearning [68], or quantization [33, 32]. The third category, backdoor sample detection, focuses on identifying triggered inputs at inference time by analyzing features like internal activations [15] or prediction consistency under input transformations [38, 16].
@@
Hardware Assisted Attacks. Hardware runtime attacks exploit physical vulnerabilities to disrupt inference without modifying the hardware design itself. Techniques include Rowhammer-induced bit flips and fault injection, which corrupt memory or logic values at runtime [39, 24, 37]. Such attacks are fundamentally different from HAMLOCK. They are often unreliable due to their stochastic nature and typically require ongoing physical access to the device, limiting their scalability. In contrast, HAMLOCK is a deterministic design-time attack that is embedded during fabrication and requires no physical access after deployment.
Hardware design-time attacks are also distinct. They primarily focus on inducing misclassifications on clean inputs (i.e., triggerless attacks) [34, 10, 67], whereas HAMLOCK is a trigger-based backdoor attack. Furthermore, these attacks often require modifying large functional units like MAC arrays [34] or ReLU trees [11], resulting in significant hardware overhead. Thanks to our hardware-model co-design, HAMLOCK requires only a few simple comparator units, leading to an effective attack with much smaller hardware footprint.

### v2

@@ 4 Experiment Setup
We describe experimental setups in Section˜4.1 and then introduce the model-level defenses for evaluation in Section˜4.2.
@@ 4.1 Experimental Setup
Backdoor related settings. The backdoor triggers on these images are $3\times 3$ squares whose values are either fixed (single neuron weight optimization and multi-neuron attacks) or optimized based on Eq. (2) (trigger optimization attack). These patch-based triggers are the most primitive and easy to detect triggers for model-level attacks, as demonstrated in prior works [65, 15, 23], and we will then show that hardware model co-attack can make these highly detectable triggers highly evasive. Some illustrative examples for different datasets are shown in Figure˜5 in the Appendix. For our multi-neuron attack, we randomly select 3 layers, each with 1 trigger neuron, constituting a 3-neuron attack.
@@
For model-level detections, we report the number of times our backdoored model is flagged as malicious, consistent with prior work [4]. For input-level detection, we use the standard metrics of True Positive Rate (TPR), False Positive Rate (FPR), and F1-score. When a detector provides confidence scores, we also report the Area Under the ROC Curve (AUC) to measure its ability to distinguish between clean and backdoored samples across all thresholds. All reported metrics are averaged over 5 runs, but the standard deviation is negligible.
@@ 4.2 Model-level Defenses
We select both classical and state-of-the-art model-level defenses to broadly assess the stealth of HAMLOCK without adaptive optimization. Our goal is not to propose one primitive attack that defeats all future defenses, instead, we aim to initiate investigation on the elevated threat from this co-attack paradigm. These defenses are grouped into four categories: (1) Backdoor model detection: Neural Cleanse [65], MNTD [69]; (2) White-box backdoor sample detection: IBD-PSC [23] and TED [42]; (3) Black-box backdoor sample detection: STRIP [15] and BBCAL [24]; and (4) Backdoor mitigation: Finetuning [51], FinePruning [37] and BEAGLE [9]. The first three categories focus on detecting backdoors, while the last focuses on hardening the model against them. Our attack is a post-training supply-chain attack [4] and hence, is not compatible with during-training defenses [60, 19].
@@ 5.1 Effectiveness without Defense
First, it maintains a high clean accuracy. Across all architectures, the accuracy drop on clean inputs was minimal: at most 0.3% for MNIST, 1.7% for CIFAR-10, 1.0% for GTSRB, and 2.6% for ImageNet. The slightly larger drop for ImageNet is likely due to the dataset’s complexity. Second, the model’s attack success rate is effectively zero. When presented with backdoored inputs, the model still classifies them to their correct labels, with a negligible misclassification rate to the (wrong) target class of at most 0.6%. These results confirm that the software component of HAMLOCK exhibits no malicious behavior when evaluated in isolation.
However, once the model is deployed on the Trojan-infected hardware, the attack becomes fully active. The hardware Trojan detects the trigger and activates its payload, causing the ASR to jump to 100% across all datasets and architectures for the 1-N attack. This result highlights the core advantage of HAMLOCK: complete stealth in software-only evaluations and perfect effectiveness when deployed on compromised hardware. The 3-N attack is slightly inferior to the 1-N attack because it lacks direct access to the trigger regions for effective optimization. Instead, it relies on indirect optimization from randomly sampled deeper layers.
Table 2: Ablation studies of HAMLOCK on VGG-16 (CIFAR-10). We report ASR to measure attack effectiveness and clean accuracy to measure utility.
(a) Impact of calibration set size on attack effectiveness
Calib. size 5 10 25 50 75 100 150 200 300 400 500
ASR (%) 63.7 70.8 74.2 82.2 92.4 93.6 93.6 95.5 96.1 96.1 96.7
Clean Acc. (%) 89.1 89.5 91.6 91.3 90.6 91.1 91.4 91.8 92.1 91.7 91.8
(b) Impact of calibration set containing samples from only a single CIFAR-10 class
Chosen Class 0 1 2 3 4 5 6 7 8 9
ASR (%) 95.0 94.0 94.0 90.0 92.0 93.0 91.0 95.0 92.0 91.0
Clean Acc. (%) 91.2 91.1 90.9 91.2 90.9 91.1 91.0 91.1 91.2
(c) Impact of neuron injection strategy and total injected neurons
# of Neurons 1 2 3 4 5 6 7 8 9
Density (3 layers) 74.0 90.1 93.6 93.9 95.2 96.0 95.9 96.1 96.3
Clean Acc. (%) 92.3 91.7 91.1 90.6 90.2 89.8 89.2 88.9 88.5
Depth (1/layer) 74.0 90.1 93.6 94.7 95.7 96.4 96.4 97.3 97.7
Clean Acc. (%) 92.3 91.7 91.1 90.8 90.4 89.9 89.8 89.5 89.3
Ablation on multi-neuron attacks. Our multi-neuron attack relies on monitoring neurons distributed over multiple, randomly selected layers. The default configuration selects $r=3$ random layers with $k=1$ trigger neuron in each layer (3 total neurons). We analyze how the two primary hyperparameters $r$ and $k$ of the multi-neuron attack impact the overall attack effectiveness, measured in ASR.
First, we measure the impact of the calibration set size with other parameters fixed in 2(a). The results show that HAMLOCK achieves reasonable effectiveness even with minimal data, reaching 70% ASR with only five calibration samples. The effectiveness steadily increases up to 98% ASR when using 500 samples. Secondly, the distribution of the calibration set, with other parameters fixed, does not significantly impact effectiveness, as shown in Table 2(b). High effectiveness (ASR >90%) is maintained even when the calibration set consists exclusively of samples from a single, particular class. This confirms the robustness of the activation separation optimization in Eq. (5). On the other hand, clean accuracy remains almost constant with calibration set using samples from a particular class.
Finally, we measure the impact of the number of trigger neurons on ASR. We test two strategies: 1) keeping the neurons fixed within 3 layers (our default layer count), and 2) distributing the neurons across a larger number of randomly selected layers, with 1 neuron in each layer. For 1), given a fixed total number of neurons (e.g., 5 neurons), we sequentially assign one trigger neuron slot per layer, starting from the earliest random layer, until all trigger neurons are assigned. In the 5-neuron case, this results in 2 trigger neurons in the first two layers and 1 neuron in the third layer. The results, shown in Table 2(c), indicate that as the number of neurons increases, the attack effectiveness steadily improves from 74% ASR for the 1-neuron case to 97.7% ASR for the 9-neuron case. Furthermore, as the number of neurons increases, the clean gradually decreases from 92.3% (for the 1-neuron case) to 89.3% (for the 9-neuron case). Notably, we find that distributing 1 neuron over more layers is slightly more effective than clustering more neurons within fewer layers.
In conclusion, the ablation study confirms that our default setting of 3 distributed trigger neurons with 100 calibration data is highly effective (94% ASR) with negligible hardware overhead and clean accuracy drop.
@@ 5.2 Evading White-box Defenses
This failure stems from a violation of each defense’s core assumptions. NC is ineffective because it searches for a small trigger pattern that causes misclassification. Our backdoored model, despite using a trigger that is very easy to reverse engineer with NC, it never misclassifies triggered inputs on its own, leaving NC with no trigger-label connection to find. Similarly, MNTD fails because it looks for statistical anomalies in neuron weights, but our attack’s minimal modification to few neurons does not create a detectable anomaly. This demonstrates that defenses examining a model’s behavior or structure are ill-equipped for a co-design attack where the malicious logic remains dormant in the software.
@@
Backdoor sample detection. We next evaluate our models against defenses that detect backdoor samples, and results are in Table˜3. Across all experimental settings, our three attack variants consistently evade detection. This is evidenced by uniformly low TPR (at most 12.4%), FPR (at most 11.5%), and F1 scores (at most 0.2). The proximity between TPR and FPR, along with AUC scores near 0.5, confirms that these defenses perform no better than random guessing at distinguishing our backdoored samples from clean ones.
@@
Effectiveness under lightweight retraining. Finally, we evaluate HAMLOCK against retraining defenses like fine-tuning and fine-pruning, with results presented in Table˜6. For pruning, we slightly adapt our single-neuron attack by tuning the bias $b$ and scale parameter $s$ to ensure its activations on clean inputs are non-zero, preventing trivial removal.
Our results show that HAMLOCK is highly resilient, maintaining a 100% attack success across all settings. Notably, this includes resilience against the advanced Beagle fine-tuning strategy [9], even under a worst-case assumption where the defender has access to backdoor samples generated from our exact attack. This resilience is twofold. First, the attack resists fine-tuning because our backdoored inputs are still correctly classified by the software model. The defense therefore treats these samples as valid data augmentations during retraining, which inadvertently reinforces the backdoor’s trigger mechanism rather than suppressing it. Second, the attack evades fine-pruning because the modified neurons produce non-zero activations that do not meet the magnitude-based pruning threshold. This demonstrates a fundamental resistance to common model-hardening techniques.
@@ 5.3 Evading Black-box Defenses
For completeness, we also evaluate black-box defenses that only require input-output access to the model. We test the robustness of both HAMLOCK and DFBA attacks against two defenses: the classic STRIP [15] and the more recent BBCAL [24]. Crucially, for HAMLOCK, we evaluate two distinct, realistic scenarios: 1) the dormant software model, as would be inspected by a Model Zoo maintainer before deployment, and 2) the active, hardware-hosted model, as would be tested by an end-user. The latter case is unique because the hardware enables misclassifications on backdoored inputs, which the software-only model does not produce.
The results of MNIST, GTSRB, CIFAR10, and Imagenet are presented in Table˜4 and Table˜5. Our HAMLOCK attack successfully bypasses both defenses in both scenarios. While BBCAL sometimes shows a high True Positive Rate (TPR), this comes at the cost of an equally high False Positive Rate (FPR), resulting in AUC scores near 0.5—no better than random guessing. The baseline DFBA attack is also largely evasive, with one minor exception: on ImageNet, STRIP achieves a 0.63 AUC score, which is still considered ineffective for reliable detection.
@@ 5.4.1 Negligible Hardware Overhead
The hardware footprint is usually measured by additional area and power added to the overall circuit. A significant area and/or power overhead can be detected using side-channel analysis [58, 29, 5], while small overheads will simply be masked by the variation of the hardware costs during the fabrication process, making the detection of these footprints almost impossible. Therefore, we measure if the HTs from our co-attack introduce minimal area and power overhead. Table˜7 summarizes the results for the three model architectures considered in this paper: LeNet, VGG-16 and ResNet-18. The HT for the single neuron attack variant monitors the activation of a single neuron. The area overhead is capped at 0.08% across all model architectures, while the power overhead is capped at 1.14% for the VGG-16 model, and the rest simply drop to 0.02%. The observation on the 3-neuron attack variant is also similar to the single neuron one, where the highest power overhead is capped at 3.4% for VGG16, while the rest are similarly below 0.05%. The area overhead is negligible with the maximum overhead being 0.1%.
@@ 5.4.2 Diverse Trigger Conditions
Our hardware-based backdoor approach allows us to compose multiple, simple triggers into diverse and sophisticated activation conditions, a capability far exceeding that of software-only attacks. We demonstrate three such compositions: combinational, sequential, and temporal triggers, all can be implemented with negligible hardware overhead.
A combinational trigger uses simple hardware logic (e.g., AND/OR gates) to combine the outputs of multiple, independent trigger detectors. For example, an AND gate requires multiple conditions to be met simultaneously (e.g., a specific road sign and foggy weather in an autonomous driving scenario), while an OR gate allows any one of several conditions to activate the payload. Such a strategy allows the attacker to split a complex trigger into smaller, stealthier pieces that are only malicious when they co-occur. We perform a preliminary experiment using two individual triggers, $3\times 3$ triggers at the bottom right and bottom left corners and each of them individually trigger a unique neuron, with the weight optimization method in Section˜3.2.1. Table˜8 shows the results, where the AND and OR logic are implemented with negligible overhead.
Beyond simple combinations, the hardware enables finite state machine (FSM) based complex sequential triggers. A sequential trigger activates the payload only after multiple, distinct trigger patterns are observed in a specific order. A temporal trigger, implemented with a simple counter circuit, activates only after a set duration has passed or some number of inferences have been made. This allows for long-term dormant attacks, such as a backdoor in an autonomous vehicle that only manifests after a specific mileage, making the resulting failure indistinguishable from natural system degradation.
@@ 6 Related Work
Model-level backdoor attacks. Existing backdoor attacks include data poisoning [17, 61, 49, 7], model parameter modification [4, 21], and model architecture modification [3, 56] approaches. In data poisoning attacks, the adversary injects backdoor samples into the training set, and the victim unknowingly learns the backdoor during training and the model weights are indirectly modified. Other backdoor attacks alter the model weights directly [21, 4] or make architectural changes [3, 47]. Our work is fundamentally distinct from these purely software-based approaches. The HAMLOCK software model contains no functional backdoor path and does not cause misclassifications on its own, making it significantly stealthier, as demonstrated in Section˜5.2.
@@
Model-level backdoor defenses. Existing model-only defenses include during-training and post-training defenses. During-training defenses require filtering our bad training data [60, 28, 35] or suppress negative impact from the bad data [67, 57, 25]. Since our attack is a supply-chain attack [4, 17] introduced post training, these defenses do not apply. Post-training defenses, evaluated in Section˜5.2, are relevant to our attack and can be grouped into three categories. First are backdoored model detection methods, which aim to identify if a model has been compromised [65, 66, 68]. Second, model hardening defenses attempt to remove the backdoor’s effect through techniques like lightweight retraining [51, 37], adversarial unlearning [72], or quantization [33, 32]. The third category, backdoor sample detection, focuses on identifying triggered inputs at inference time by analyzing features like internal activations [14] or prediction consistency under input transformations [39, 15].
@@
Hardware runtime attacks. Hardware runtime attacks exploit physical vulnerabilities to disrupt inference without modifying the hardware design itself. Techniques include Rowhammer-induced bit flips and fault injection, which corrupt memory or logic values at runtime [40, 22, 38]. Such attacks are fundamentally different from HAMLOCK. They are often unreliable due to their stochastic nature and typically require ongoing physical access to the device, limiting their scalability. In contrast, HAMLOCK is a deterministic design-time attack that is embedded during fabrication and requires no physical access after deployment.
Hardware design-time attacks. These attacks relate to malicious modification of NN hardware at design stage. We provide detailed comparisons below.
Pattern matching HTs: Several previous work leverage pattern matching HTs on the input data or network activations to detect triggers [71, 73, 10]. Ye et al. [71] checks the bit representation of a small input patch (e.g., matching a 72-bit input from a 3×3 RGB window) of a clean image to decide on the input to be misclassified. While this approach might be extended to the backdoor setting by checking the bit patterns of the backdoor trigger region, it is unknown whether clean and backdoor samples can be sufficiently separated using such exact pattern match. More importantly, the hardware complexity for this method linearly increases with trigger size or the number of trigger patterns, making it unscalable.
Other existing works require more complex checks to match the patterns of the specific intended input to misclassify. The HTA [10] checks the internal activation vector of a small percentage of neurons (e.g., 0.03%), while the MTA attack [73] correlates statistical features, such as white/black ratios and memory access patterns. When adapting these attacks to backdoor settings, diverse inputs containing the same trigger pattern can induce different statistical features. Therefore, successfully generalizing these checks for backdoor attacks requires significant hardware overhead in the trigger circuit. This demands complex, sequential units, e.g., a control circuit implemented by an FSM to track and correlate internal states, and arrays of multi-bit digital comparators to perform the multi-dimensional vector similarity check. Moreover, these attacks require an additional payload circuit to physically inject the misclassification perturbation, compounding the area and power overhead.
Model architecture HTs: Other works, such as the Hu-Fu attack [34], train a model to misclassify a particular clean input only when a dedicated subnet of parameters is physically activated. This mask for activation is often realized using a physical wire, requiring an attacker to physically access the deployed hardware at inference time. While potentially adaptable to backdoor settings, this attack must dedicate a significant portion of the model architecture for the malicious function (e.g., reserving 5 cross-shaped weights out of 9 in a 3×3 kernel for a ResNet-20 model), which effectively increases the overall hardware overhead. Crucially, relying on direct physical access via an external wire makes this attack largely infeasible against remotely deployed hardware.
Benefits of HAMLOCK: HAMLOCK addresses the limitations above by optimizing the model weights to converge the diverse trigger-related activations into a uniquely high activation on a few ($\leq 3$) neurons (Section˜3.2). This co-design efficiently reduces the demanding hardware task from a complex vector match or a subnetwork to a simple threshold check on a few activation bits (e.g., the most significant bit), providing a ultra-low-overhead and scalable strategy. Furthermore, this extreme reduction in essential hardware logic leaves room for implementing additional temporal and spatial trigger logic, which allows for wide variety of payloads, e.g., sequential misclassifications for distributed trigger patterns, time-based, or random misclassifications temporally (Section˜5.4.2). Regarding the attack payload, many existing HTs focus on simpler class-agnostic or untargeted misclassifications with unpredictable output class, while HAMLOCK enables class-agnostic and class-specific targeted misclassifications (Section˜3.3).